OpenStudy (anonymous):

Please help!!!!!!! fan and medals!!!!! Which relation is a function? A. {(1, 2); (2, 3); (3, 4); (4, 5)} B. {(1, 2); (3, 2); (3, 3); (4, 2)} C. {(1, 2); (2, 3); (3, 4); (2, 5)} D. {(1, 2); (1, 3); (1, 4); (1, 5)}

OpenStudy (anonymous):

@mathstudent55 @joshhfers

OpenStudy (anonymous):

Plot the points for each of the four answers on four separate graphs. If you can draw a vertical line though any two or more points, then it is not a function.

OpenStudy (anonymous):

i got thanks

OpenStudy (anonymous):

can you help with more

OpenStudy (anonymous):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 10 when x = 2. A. y = 5x B. y = x – 8 C. y = x + 8 D. y = 2x + 6

OpenStudy (anonymous):

If you look up direct variation. It states y=Kx Where K is a constant. So it looks like A is the correct answer.